the length of the garden should be 2 less than 3 times the width. The area should be 16 cm. Find the dimensions & perimeter of the garden.​
Answer by greenestamps(13216)



x = width
3x-2=length

The area (length times width) is 16:

x(3x-2)=16
3x^2-2x=16
3x^2-2x-16=0
(3x-8)(x+2)=0

x=8/3 or x=-2

Reject the negative solution, since it makes no sense for the actual problem.

ANSWERS:
width = x = 8/3
length = 3x-2 = 8-2=6
area = (8/3)(6) = 16
perimeter = twice width plus twice length = 2(8/3)+2(6) = 16/3+12 = 52/3 or 17 1/3